# Scaling Laws for Uncertainty in Deep Learning

### v1

@@ 4.1.2 Vision Transformer
We explore uncertainty scaling trends using Vision Transformer (ViT) architectures (Dosovitskiy et al., 2021). We use a compact ViT model trained on both CIFAR-10 and ImageNet32. The architecture consists of a $6$-layer transformer with $8$ attention heads per layer and input images are tokenized in $4$ patches. We apply mc dropout with a fixed rate of $0.5$ both on the embeddings and in the transformer blocks. We conduct an ablation study in Fig. 5 to highlight how different experimental settings lead to completely different uncertainty behaviors.
It is interesting to observe the effect of the cosine annealing learning schedule, also visible in Fig. 5 when training on CIFAR-10, suggesting that early-phase uncertainty dynamics in transformers can be sensitive to optimization strategies. Results on ImageNet32 are reported in Appendix B .
@@ 4.2 Text classification
Figure 4: Algorithmic dataset: GPT-2 model on the Algorithmic Dataset to learn the modulo addition operation. Results obtained by training for 10 0001000010\,00010 000 epochs averaged over $3$ folds.
We further investigate the language modality by assessing uncertainty estimates on the pre-trained Phi-2 model (Abdin et al., 2024), applying a Laplace approximation to the posterior over the LoRA parameters (Yang et al., 2024). We fine-tuned the model on qqp and ARC datasets observing that the uncertainties remained flat for every data subset used for fine tuning (see Fig. 15). This saturation effect is likely due to the massive dataset used for pre-training, limiting the expressiveness of the model uncertainty on comparatively much smaller data for fine tuning.
@@ 4.2.1 GPT-2 on Algorithmic Dataset
We train a GPT-2 model to solve modular arithmetic problems from a synthetic dataset based on the setup in Power et al. (2022). The model learns to predict the token following the equals sign. We only experiment with mc dropout with rate $0.1$ to estimate predictive uncertainty. Interestingly, more evident uncertainty scaling patterns only emerge after extensive training, suggesting potential links between grokking dynamics (Belkin et al., 2019) and uncertainty behavior. In Fig. 4 we show predictive uncertainties when training the model on increasing percentages (from $5\%$ to $90\%$) of the MODULO 97 algorithmic dataset.
Figure 5: ViT-small uncertainty training dynamics on CIFAR-10 dataset: In (a) we use mc dropout ($p=0.5$) and we train the model for $500$ epochs with Adam optimizer (Kingma and Ba, 2017) and cosine annealing. In (b) we train the same model for $2500$ epochs and fixed learning rate $10^{-4}$. The training dynamics show that the the shape/speed of convergence of of tu, au, eu strongly depends on the optimization trajectories underneath.
Figure 6: ResNets on CIFAR10-C dataset: mc dropout with $p=0.2$ in (a) and $p=0.5$ in (b). For models trained on incrementally larger training subsets of CIFAR-10, we report the predictive uncertainties when testing on the (whole) CIFAR10-C dataset, averaged over all corruption levels ($1$-$5$) and corruption types considered.
Figure 7: Uncertainty scaling with model parameters: An extensive range of ResNets architectures on CIFAR-10 dataset - mc dropout ($p=0.5$). We train all these architectures for $400$ epochs and fixed learning rate $10^{-3}$.
Figure 8: Uncertainty estimates for ResNets on ImageNet-32 dataset: The first two subplots report mc dropout ($p=0.5$) uncertainties for ResNet-18 and ResNet-34, trained with fixed learning rate $10^{-3}$ on $9$ increasing subsets of the training data. We also report uncertainties obtained through mcmc considering only the first layer stochastic (mcmc (1)) and only the first two layers stochastic (mcmc (2)) for ResNet-18 trained on $4$ increasing subsets of the training data.
@@ 5 Theoretical connections
The theory of identifiable models has extensively analyzed asymptotic behaviors, such as posterior contraction and convergence of test loss and generalization error. In this section, we recall how predictive uncertainty contracts with increasing data in Bayesian linear regression, highlight its connection to Watanabe’s Generalization Error, and outline links to Singular Learning Theory.
5.1 tu scaling in Identifiable Parametric Models
We consider Bayesian linear regression $y={{\boldsymbol{\mathbf{\theta}}}}^{\top}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}})+\epsilon$, where ${{\boldsymbol{\mathbf{\theta}}}}\in\mathbb{R}^{P}$ are parameters of interest, and ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}})=[\phi_{1}({{\boldsymbol{\mathbf{x}}}}),\phi_{2}({{\boldsymbol{\mathbf{x}}}}),\dotsc,\phi_{P}({{\boldsymbol{\mathbf{x}}}})]^{\top}$ are the basis functions, and assume zero-mean noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We define the likelihood over $N$ iid observations $\{{{\boldsymbol{\mathbf{X}}}},{{\boldsymbol{\mathbf{y}}}}\}=\{({{\boldsymbol{\mathbf{x}}}}_{i},y_{i})\}_{i=1}^{N}$,
@@
By assuming a conjugate Gaussian prior $p({{\boldsymbol{\mathbf{\theta}}}})=\mathcal{N}({{\boldsymbol{\mathbf{\theta}}}}|{{\boldsymbol{\mathbf{m}}}}_{0},{{\boldsymbol{\mathbf{S}}}}_{0})$ the posterior is also Gaussian with mean ${{\boldsymbol{\mathbf{m}}}}_{N}={{\boldsymbol{\mathbf{S}}}}_{N}\left({{\boldsymbol{\mathbf{S}}}}_{0}^{-1}{{\boldsymbol{\mathbf{m}}}}_{0}+\sigma^{-2}{{\boldsymbol{\mathbf{\Phi}}}}^{\top}{{\boldsymbol{\mathbf{y}}}}\right)$ and covariance ${{\boldsymbol{\mathbf{S}}}}_{N}=\left(\sigma^{-2}{{\boldsymbol{\mathbf{\Phi}}}}^{\top}{{\boldsymbol{\mathbf{\Phi}}}}+{{\boldsymbol{\mathbf{S}}}}_{0}^{-1}\right)^{-1}$. The predictive posterior for a new test points $({{\boldsymbol{\mathbf{x}}}}_{*},y_{*})$ is also Gaussian, $p(y_{*}|{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}},{{\boldsymbol{\mathbf{X}}}})=\mathcal{N}\Big{(}y_{*}|{{\boldsymbol{\mathbf{m}}}}_{N}^{\top}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*}),\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}}_{*})\Big{)}$ with $\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}}_{*})=\sigma^{2}+{{\boldsymbol{\mathbf{\phi}}}}(x_{*})^{\top}{{\boldsymbol{\mathbf{S}}}}_{N}{{\boldsymbol{\mathbf{\phi}}}}(x_{*}).$
The predictive variance $\mathrm{Var}[y_{*}|{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}},{{\boldsymbol{\mathbf{X}}}}]$ decomposes into the uncertainty of data $\sigma^{2}$ (au), and the uncertainty of parameters ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})^{\top}{{\boldsymbol{\mathbf{S}}}}_{N}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})$ (eu). We stress that eu is made available only by taking the Bayesian approach of treating ${{\boldsymbol{\mathbf{\theta}}}}$ as a random variable. It can be shown that $\sigma^{2}_{N+1}({{\boldsymbol{\mathbf{x}}}}_{*})\leq\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}}_{*})$ – the posterior distribution becomes narrower as additional data points are observed (Qazaz et al., 1997). Moreover, in the limit of $N\rightarrow\infty$ we get $\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})\rightarrow\sigma^{2}$: in the infinite data regime, the predictive uncertainty converges to its irreducible au component.
@@ 5.2 Singular Learning Theory
Our speculative theoretical link is slt, which provides insights into the learning dynamics of highly overparameterized models. The Fisher information of deep neural networks is often singular at certain parameter values; despite forming a measure-zero subset, these singularities significantly influence learning. slt proves that the asymptotic properties of learning are shaped by the geometry near such degenerate points (Watanabe, 2009, 2018). One main quantity of interest in this framework is the generalization error $G$. Here we limit ourself to connect this to the tu in Bayesian Linear Regression. Using Watanabe’s notation
[…]
@@
is the negative marginal log likelihood (NMLL) of the data ${{\boldsymbol{\mathbf{X}}}}_{N}$ of size $N$. The generalization error $G_{N}=\mathbb{E}_{p(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},\hat{{{\boldsymbol{\mathbf{\theta}}}}})}[F_{N+1}]-F_{N}$ measures the expected increase in marginal likelihood when training with one additional data point (See Section A.2) and it can be rewritten as:
$G_{N}=-\mathbb{E}_{\underbrace{\scriptstyle p(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},\boldsymbol{\theta})}_{p(y)}}\Big{[}\log\underbrace{p\big{(}y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},{{\boldsymbol{\mathbf{X}}}}_{N},{{\boldsymbol{\mathbf{y}}}}_{N}\big{)}}_{q_{N}(y)}\Big{]},$ (8)
@@
which reduces to the log posterior predictive distribution for the $(N+1)$’th datapoint given the first $N$ datapoints, under the true process
$y_{N+1}\sim\mathcal{N}\Big{(}y_{N+1}|{{\boldsymbol{\mathbf{\theta}}}}_{\mathrm{true}}^{\top}{{\boldsymbol{\mathbf{x}}}}_{N+1},\sigma_{\mathrm{true}}^{2}\Big{)}.$ (9)
By denoting the true process by $p(y)$ and the posterior predictive from $N$ data by $q_{N}(y)$ we can manipulate the expression of the generalization error to obtain:
@@
$G_{N}=-\underbrace{\frac{1}{2}\log{(2\pi e\sigma_{\mathrm{true}}^{2})}}_{\text{aleatoric uncertainty}}-\underbrace{\mathrm{KL}\big{[}p(y)||q_{N}(y)\big{]}}_{\text{epistemic uncertainty}},$ (10)
@@
where $\mathrm{KL}[p(y)||q_{N}(y)]$ is the kl divergence between the true and the posterior predictive distribution, quantifying the epistemic uncertainty arising from limited knowledge of the true model parameters. By looking at the asymptotic expression of tu as $N\rightarrow\infty$ and by taking, without loss of generality, identity basis functions, we get:
$\mathrm{TU}({{\boldsymbol{\mathbf{x}}}}_{N+1})$ $=\mathbb{H}[q_{N}(y)]$ (11)
@@
$=\frac{1}{2}\log(2\pi e\sigma_{\mathrm{true}}^{2})+\frac{1}{2(N+1)}{{\boldsymbol{\mathbf{x}}}}_{N+1}^{\top}{{\boldsymbol{\mathbf{\Sigma}}}}_{X_{N+1}}^{-1}{{\boldsymbol{\mathbf{x}}}}_{N+1}+O\left(\frac{1}{(N+1)^{2}}\right).$ (12)
@@
As more data is collected, the predictive posterior $q_{N}(y)$ converges toward the true predictive distribution $p(y)$. Consequently, the generalization error $G_{N}$ asymptotically approaches the irreducible aleatoric uncertainty. Similarly, the tu converges to the aleatoric component as the epistemic part vanishes. In Watanabe (1999) it’s proved an asymptotic expansion of $G_{N}$ which is proportional to the effective dimensionality of the model.
@@
We hypothesize that the theoretical tools provided by this framework, closely linked to concepts from Statistical Physics, may offer a promising foundation to explain some of the scaling behaviors we have observed. In particular, we find recent efforts to computationally characterize the effective dimensionality of deep learning models especially compelling (Lau et al., 2024; Chen et al., 2024), and we aim to investigate formal connections in future extensions of this work.

### v2

@@ 4.1.2 Vision Transformer
We explore uncertainty scaling trends using Vision Transformer (ViT) architectures (Dosovitskiy et al., 2021). We conduct a study to highlight how different experimental settings lead to different uncertainty behaviors (Figure˜8). It is interesting to observe the effect of annealing the learning rate with a cosine scheduler when training on CIFAR-10, suggesting that early-phase uncertainty dynamics in transformers can be sensitive to optimization strategies. Results on ImageNet32 are reported in Figure˜13.
ViT uncertainty scaling with model parameters
In ViT models, eu decreases with model size (negative $\gamma$ across all patch sizes and depths), as it can be observed in Figure˜9. We believe that larger ViTs produce smoother representations and show stronger inductive bias toward stable features; consequently, different dropout samples yield increasingly similar predictive distributions. In this setting, model size reduces ensemble diversity, and therefore eu contracts.
Figure 9: ViT uncertainty scaling with model parameters: We train different sizes of ViT by varying the patch size (88, $16$), the number of layers (66, $12$) and the number of attention heads (each point corresponds to $8$, $16$ and $24$ attention heads respectively). We use mcd (with fixed $p=0.5$ both in the embeddings and in the transformer blocks). au shows minimal dependence on $P$. tu mirrors the trend of the epistemic component. Overall, the slopes are small in magnitude ($|\gamma|<0.2$), indicating that uncertainty decreases only slowly with model size.
@@ 4.2 Text classification
We firstly investigate the language modality by assessing uncertainty estimates on the pre-trained Phi-2 model (Abdin et al., 2024), applying a Laplace approximation to the posterior over the LoRA parameters (Yang et al., 2024). We fine-tuned the model on qqp and ARC datasets observing that the uncertainties remain flat for every data subset used for fine tuning (see Figure˜18). This saturation effect is likely due to the massive amount of data used for pre-training, limiting the expressiveness of the model uncertainty on comparatively much smaller data for fine-tuning.
@@ 4.2.1 GPT-2 on Algorithmic Dataset
Figure 10: Algorithmic dataset: GPT-2 model on the Algorithmic dataset to learn the modulo addition operation. Results are obtained by training for 10 00010\,000 epochs and averaging over 33 seeds.
We train a GPT-2 model to solve modular arithmetic problems of the form $(a+b)\bmod m=$, where $m$ is the modulus, $a$ and $b$ are integers from 0 to $m$, and the model learns to predict the token after the == sign. This setup and dataset differ significantly from the vision tasks considered so far, thus providing a complementary testbed for our hypotheses. We experiment using mcd with a dropout rate of $p=0.1$. In Figure˜10 we show predictive uncertainties when training the model on increasing percentages (from $5\%$ to $90\%$) of the $m=97$ algorithmic dataset. Similarly to previous experiments, we observe clear scaling laws for all uncertainty metrics considered. Interestingly, more evident uncertainty scaling patterns only emerge after extensive training, suggesting potential links between grokking (Belkin et al., 2019) and uncertainty behavior. The slight uncertainty increase in the last $N$ subset of Figure˜10 likely stems from convergence dynamics and limitations of mcd.
4.3 Extrapolation on a toy dataset
Figure 11: Uncertainty scaling of an mlp on the two_moons dataset: We use Bayesian logistic regression with normal priors on an mlp with two hidden layers of size $32$ and $\tanh$ activation functions. Posterior inference is performed with hmc sampling ($200$ posterior samples after 1,0001{,}000 warm-up steps). Training is carried out on $12$ log-spaced values ranging from $100$ to $1M$ and 44 random seeds. Uncertainties are computed on the held-out 5,0005{,}000 test set.
We conduct an experiment using a simple mlp on the two_moons  [^1] Dataset available through Scikit-learn (Pedregosa et al., 2011) toy classification dataset to demonstrate the practical application of uncertainty scaling laws. As illustrated in Figure˜11, the eu exhibits a clear power-law decay across the observed range of $N$. As expected, the au contribution settles down and plateaus for large values of $N$, representing the irreducible noise inherent in the task. The tu curve eventually diverges from the power-law trend as it approaches this aleatoric floor. Beyond the extrapolation boundary, the predicted eu remains highly consistent with the actual observed scaling. This is particularly powerful because it demonstrates that training on a relatively small regime, up to just a few thousand data points, allows us to accurately predict the uncertainty behavior we can expect when scaling the data to the order of millions of samples.
@@ 5 Theoretical connections
The theory of identifiable models has extensively analyzed asymptotic behaviors, such as posterior contraction and convergence of test loss and generalization error. In this section, we recall how predictive uncertainty contracts with increasing data in Bayesian linear regression, highlight its connection to Watanabe’s Generalization Error, and outline links to slt.
5.1 Total Uncertainty scaling in Identifiable Parametric Models
We consider Bayesian linear regression $y={{\boldsymbol{\mathbf{\theta}}}}^{\top}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}})+\epsilon$, where ${{\boldsymbol{\mathbf{\theta}}}}\in\mathbb{R}^{P}$ are parameters of interest, and ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}})=[\phi_{1}({{\boldsymbol{\mathbf{x}}}}),\phi_{2}({{\boldsymbol{\mathbf{x}}}}),\dotsc,\phi_{P}({{\boldsymbol{\mathbf{x}}}})]^{\top}$ are basis functions, and assume zero-mean noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We define the likelihood over $N$ iid observations $\{{{\boldsymbol{\mathbf{X}}}},{{\boldsymbol{\mathbf{y}}}}\}=\{({{\boldsymbol{\mathbf{x}}}}_{i},y_{i})\}_{i=1}^{N}$,
@@ 5.1 Total Uncertainty scaling in Identifiable Parametric Models
By assuming a conjugate Gaussian prior $p({{\boldsymbol{\mathbf{\theta}}}})=\mathcal{N}({{\boldsymbol{\mathbf{\theta}}}}|{{\boldsymbol{\mathbf{m}}}}_{0},{{\boldsymbol{\mathbf{S}}}}_{0})$ the posterior is also Gaussian with mean ${{\boldsymbol{\mathbf{m}}}}_{N}={{\boldsymbol{\mathbf{S}}}}_{N}\left({{\boldsymbol{\mathbf{S}}}}_{0}^{-1}{{\boldsymbol{\mathbf{m}}}}_{0}+\sigma^{-2}{{\boldsymbol{\mathbf{\Phi}}}}^{\top}{{\boldsymbol{\mathbf{y}}}}\right)$ and covariance ${{\boldsymbol{\mathbf{S}}}}_{N}=\left(\sigma^{-2}{{\boldsymbol{\mathbf{\Phi}}}}^{\top}{{\boldsymbol{\mathbf{\Phi}}}}+{{\boldsymbol{\mathbf{S}}}}_{0}^{-1}\right)^{-1}$.
The predictive posterior for a new test points $({{\boldsymbol{\mathbf{x}}}}_{*},y_{*})$ is also Gaussian, $p(y_{*}|{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}},{{\boldsymbol{\mathbf{X}}}})=\mathcal{N}\Big(y_{*}|{{\boldsymbol{\mathbf{m}}}}_{N}^{\top}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*}),\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}}_{*})\Big)$ with $\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}}_{*})=\sigma^{2}+{{\boldsymbol{\mathbf{\phi}}}}(x_{*})^{\top}{{\boldsymbol{\mathbf{S}}}}_{N}{{\boldsymbol{\mathbf{\phi}}}}(x_{*}).$
The predictive variance $\mathrm{Var}\Big[y_{*}\mid{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}},{{\boldsymbol{\mathbf{X}}}}\Big]$ decomposes into the uncertainty of data $\sigma^{2}$ (au), and the uncertainty of parameters ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})^{\top}{{\boldsymbol{\mathbf{S}}}}_{N}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})$ (eu). It can be shown that $\sigma^{2}_{N+1}({{\boldsymbol{\mathbf{x}}}}_{*})\leq\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}}_{*})$ – the posterior distribution becomes narrower as additional data points are observed (Qazaz et al., 1997). Moreover, in the limit of $N\rightarrow\infty$ we get $\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})\rightarrow\sigma^{2}$: the predictive uncertainty converges to its irreducible au component.
@@ 5.2 Singular Learning Theory
Our speculative theoretical link is slt, which provides insights into the learning dynamics of highly overparameterized models. The Fisher information of deep neural networks is often singular at certain parameter values; despite forming a measure-zero subset, these singularities significantly influence learning. slt proves that the asymptotic properties of learning are shaped by the geometry near such degenerate points (Watanabe, 2009, 2018). Here we limit ourself to connect this to the tu in Bayesian linear regression. Using Watanabe’s notation:
$F_{N}=-\log p({{\boldsymbol{\mathbf{y}}}}_{N}|{{\boldsymbol{\mathbf{X}}}}_{N})=-\log\int p({{\boldsymbol{\mathbf{y}}}}_{N}|{{\boldsymbol{\mathbf{X}}}}_{N},{{\boldsymbol{\mathbf{\theta}}}})p({{\boldsymbol{\mathbf{\theta}}}})d{{\boldsymbol{\mathbf{\theta}}}}$ (8)
@@
is the Negative Marginal Log Likelihood (nmll) of the data ${{\boldsymbol{\mathbf{X}}}}_{N}$ of size $N$. By denoting the true conditional distribution $p(y)\equiv p(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},{{\boldsymbol{\mathbf{\theta}}}}_{\mathrm{true}})$ (e.g., $y_{N+1}\sim\mathcal{N}\big({{\boldsymbol{\mathbf{\theta}}}}_{\mathrm{true}}^{\top}{{\boldsymbol{\mathbf{x}}}}_{N+1},\sigma_{\mathrm{true}}^{2}\big)$), the generalization error $G_{N}=\mathbb{E}_{p(y)}[F_{N+1}]-F_{N}$ measures the expected increase in marginal likelihood when training with one additional data point (see Section˜A.2). This can be rewritten as:
$G_{N}=-\mathbb{E}_{p(y)}\Big[\log\underbrace{p\big(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},{{\boldsymbol{\mathbf{X}}}}_{N},{{\boldsymbol{\mathbf{y}}}}_{N}\big)}_{q_{N}(y)}\Big],$ (9)
@@
which reduces to the log posterior predictive distribution for the $(N+1)$’th datapoint given the first $N$ datapoints, under the true conditional distribution $p(y)$. By denoting the posterior predictive $q_{N}(y)$, we can manipulate the expression in Equation˜9 to obtain:
@@
$G_{N}=\underbrace{\frac{1}{2}\log{(2\pi e\sigma_{\mathrm{true}}^{2})}}_{\text{aleatoric uncertainty}}+\underbrace{\mathrm{KL}\big[p(y)||q_{N}(y)\big]}_{\text{epistemic uncertainty}},$ (10)
@@
where $\mathrm{KL}[p(y)||q_{N}(y)]$ is the kl divergence between the true and the posterior predictive distribution, quantifying the eu arising from limited knowledge of the model parameters. By looking at the asymptotic expression of tu as $N\rightarrow\infty$ and by taking, without loss of generality, identity basis functions, we get:
@@
$\begin{split}\mathrm{TU}({{\boldsymbol{\mathbf{x}}}}_{N+1})&=\mathbb{H}[q_{N}(y)]\\ &=\frac{1}{2}\log(2\pi e\sigma_{\mathrm{true}}^{2})+\frac{{{\boldsymbol{\mathbf{x}}}}_{N+1}^{\top}{{\boldsymbol{\mathbf{\Sigma}}}}_{X_{N+1}}^{-1}{{\boldsymbol{\mathbf{x}}}}_{N+1}}{2(N+1)}\\ &\quad+O\left(\frac{1}{(N+1)^{2}}\right).\end{split}$ (11)
@@
As more data is collected, the predictive posterior $q_{N}(y)$ converges to the true predictive distribution $p(y)$. Consequently, the generalization error $G_{N}$ asymptotically approaches the irreducible au. Similarly, the tu converges to the aleatoric component as the epistemic part vanishes. In Watanabe (1999) it’s proved an asymptotic expansion of $G_{N}$ which is proportional to the effective dimensionality of the model (Section˜A.2.1).
@@
Recent advances on the effective dimensionality of deep models (Lau et al., 2024; Chen et al., 2024) appear especially promising, and we intend to investigate such formal connections in future work. We further hypothesize that concepts from Statistical Physics may help explain the scaling behaviors we observed.